1. A 5 kg ball traveling to the right at 7 m/s collides with a 3 kg ball at rest. After the collision, the 3 kg ball is traveling to the right at 8.75 m/s. What is the velocity of the ball on the left after the collision?

2. A 100-kg box on the right and a 45-kg box on the left are tied together by a horizontal cable. The 100-kg box is acted upon by an external force of 560 N to the right. The frictional force exerted by the floor on the 100-kg box is 40 N and the frictional force exerted by the floor on the 45-kg box is 15 N. What is the magnitude of the tension in the connecting cable?
Answer #1

1) using conservation of momentum

m1v1i + m2v2i = m1v1f + m2v2f

5*7 + 3*0 = 5*v1f + 3*8.75

35 = 5*v1f + 26.25

v1f = 1.75 m/s ( to the right)

--------------------------------------------------

2) 560 - f1 - f2 = (m1 + m2)a ( From newton's second law)

I have assumed that f1 and f2 force act to the left because the whole system moves to right

560 - 40 - 15 = 145*a

a = 3.4827 m/s2

so

T - 15 = 45*3.4827

T = 172 N